OpenStudy (anonymous):

x^2+y^2-2x-10y+2=0 find the circles radius and center! is the answer 1,5 and sqrt 24?

hartnn (hartnn):

its absolutely correct! :) good work.

OpenStudy (anonymous):

hey do u know how to find average rate of change in function?

hartnn (hartnn):

yes do you have a specific example you need help with ?

hartnn (hartnn):

for f(x) its just \(\large \dfrac{f(x)-f(a)}{x-a}\)

OpenStudy (anonymous):

f(x)=-4x^2+2x-3 the interval is 1,2

OpenStudy (anonymous):

is the answer for dat one is 10?

hartnn (hartnn):

-35 - (-5) = -30 how did you get 10 ?

OpenStudy (anonymous):

-4(2)^2+2(2)-3-(-4(1)^2+2-3) from this

OpenStudy (anonymous):

how did you get -30?

hartnn (hartnn):

i think i did it in my mind incorrectly

OpenStudy (anonymous):

would you try doing it on paper? cuz i think i did it wrongly and i tried to find an answer from here but no one seems to reply for it

hartnn (hartnn):

yes, you're correct -4(2)^2+2(2)-3-(-4(1)^2+2-3) = - 15 - (-5) = -10 is this what u get ?

OpenStudy (anonymous):

yup

hartnn (hartnn):

then you're correct :)

OpenStudy (anonymous):

thanks

hartnn (hartnn):

welcome ^_^

OpenStudy (anonymous):

one more question! :)) do u kno anything about fog? the compostion of g and f?

hartnn (hartnn):

yes

OpenStudy (anonymous):

would you mind answering my question in a new section? that way i can give another medal

hartnn (hartnn):

as you wish :)